A ballon raising vertically with a velocity of 5m/sec drops a body from a height of 20m above the ground.The velocity of the body 1sec after release is

Grade:11

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Chakrala,

 given, 20 = 5 - 10* t2/2

or t = sqrt3

now body is drop, so

v = u + at

V = -20 + 10(t+1) = 7.32 m/s

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay